# LlavaGuard: An Open VLM-based Framework for Safeguarding Vision Datasets and Models

### v2

@@ Abstract
This paper introduces LlavaGuard, a suite of VLM-based vision safeguards that address the critical need for reliable guardrails in the era of large-scale data and models. To this end, we establish a novel open framework, describing a customizable safety taxonomy, data preprocessing, augmentation, and training setup. For teaching a VLM safeguard on safety, we further create a multimodal safety dataset with high-quality human expert annotations, where each image is labeled with a safety rating, category and rationale. We also employ advanced augmentations to support context-specific assessments. The resulting LlavaGuard models, ranging from 0.5B to 7B, serve as a versatile tool for evaluating the safety compliance of visual content against flexible policies. In comprehensive experiments, LlavaGuard outperforms both state-of-the-art safeguards and VLMs in accuracy and in flexibly handling different policies. Additionally, we demonstrate LlavaGuard’s performance in two real-world applications: large-scale dataset annotation and moderation of text-to-image models. We make our entire framework publicly available [^1] https://github.com/ml-research/LlavaGuard, including the dataset and model weights.
Safety, Safeguarding, VLM, Dataset curation, Model safeguarding
Warning: This paper contains explicit imagery, and other content that readers may find disturbing.
@@ 1 Introduction
Furthermore, emerging legal frameworks for AI, such as those in the EU (EU, 2023), US (US, 2023), and UK (UK, 2023), are pushing for advanced generative models to comply with new regulations. This has led to the proposal of various safety approaches and taxonomies to systematically assess and mitigate the risks associated with large-scale data and models (Inan et al., 2023; Wang et al., 2023; Schramowski et al., 2023; Tedeschi et al., 2024). However, prior safety research focuses primarily on the text domain, leaving a distinct lack of frameworks for the visual modality. Moreover, there is a dearth of datasets and pipelines necessary for building advanced safeguard models. Consequently, users largely have to rely on rigid NSFW classifications (Qu et al., 2024; Birhane & Prabhu, 2021; Birhane et al., 2023; Schramowski et al., 2022; NotAI-tech, 2019; Laborde, 2020), which lack the context-awareness and flexibility needed for more nuanced, fine-grained analysis.
We bridge this gap by introducing LlavaGuard (Fig. 1), a versatile framework for assessing potentially unsafe image content. LlavaGuard combines visual and textual inputs, allowing for the assessment of arbitrary safety policies to meet diverse requirements. To this end, we enhance the general capabilities of VLMs in two key ways. Firstly, building on their inherent common-sense understanding, LlavaGuard is trained with an in-depth and adaptive understanding of safety. This safety-specific training enables it to provide detailed that include an overall safety rating (safe/unsafe), a specific safety category (e.g. hate or sexual content), and a rationale that explains why the content is deemed unsafe according to the given policy. Secondly, with an advanced training setup, LlavaGuard is equipped with an abilitiy to flexibly handle a broad spectrum of policies. Given the variability in regulations—such as cannabis being illegal in some countries but legal in others—LlavaGuard can be easily adjusted to both contexts.
In summary, our contributions are as follows: (1) We establish an open framework for vision safeguards, encompassing a safety taxonomy, data preprocessing, augmentation, and training setup. (2) We construct a multimodal safety dataset with human annotations, including images labeled with a safety rating, category, and rationale (Sec. 4). (3) Based on the previous, we launch LlavaGuard, a suite of vision safeguard models based on VLMs, trained to assess visual content for safety (Sec. 5). (4) We conduct comprehensive experiments demonstrating that LlavaGuard outperforms state-of-the-art VLMs and state-of-the-art safeguards, excelling not only in accuracy but also in flexibly handling different policies (Sec. 6). (5) Finally, we validate LlavaGuard’s performance on two real-world applications: dataset annotation and moderation of generative models (Sec. 7).
@@ 2 Background
Safety Audits. Gebru et al. (2021) initiated the effort of systematically reporting visual content by advocating for meticulous documentation of datasets to promote their ethical use. Initial approaches are centered around classification tools, where common ones are convolutional (NotAI-tech, 2019; Karkkainen & Joo, 2021) and CLIP-based (Schramowski et al., 2022; Nichol et al., 2022) classifiers or human annotations (Birhane et al., 2021). In particular, NudeNet (NotAI-tech, 2019), NSFW-Nets (Falconsai, 2024; Sanali209, 2024) and Birhane & Prabhu (2021) focus on NSFW, FairFace (Karkkainen & Joo, 2021) on fairness, Q16 (Schramowski et al., 2022) on in/appropriatness, and Nichol et al. (2022) on privacy and violence.
@@
Based on these tools and efforts common large-scale datasets such as LAION (Schuhmann et al., 2022) or ImageNet (Deng et al., 2009) have undergone careful curation from different perspectives (Qu et al., 2024; Birhane & Prabhu, 2021; Birhane et al., 2023; Schramowski et al., 2022; Schuhmann et al., 2022). The resulting (unsafe) subsets serve a dual purpose. First, they are crucial in excluding content that could compromise safety during model training, ensuring a safer training environment. Second, these subsets provide valuable resources for conducting safety-oriented research. Furthermore, with the rise of models generating images, prompt testbeds such as I2P (Schramowski et al., 2023) or MAGBIG (Friedrich et al., 2024b) have been proposed for safety audits, moving beyond real images.
@@
However, the scope of these audits and auditing tools is limited by the capabilities of their underlying models, which lack the versatility and advanced common-sense understanding provided by large-scale, pre-trained VLMs. This capability is essential for effectively handling both real and synthetic images across a broad range of domains. In contrast, LlavaGuard is built on such VLMs which are not restricted to a fixed set of safety dimensions and can be easily adapted to accommodate a variety of policies.
@@
Generative AI Risk Assessment and Mitigation. In the context of generative models, most existing studies focus on the textual modality. Endeavors to systematically categorize safety risks have spurred the creation of safety taxonomies (Inan et al., 2023; Wang et al., 2023; Tedeschi et al., 2024), which provide a structured framework for assessing and mitigating risks. In particular, Inan et al. (2023) proposed a taxonomy enabling the LlamaGuard model to classify harmful prompts and responses into 6 categories. Similarly, Wang et al. (2023) proposed an 8-category taxonomy to evaluate LLMs based on different safety and trustworthiness perspectives, including robustness to adversarial attacks. These taxonomies constitute an initial stride toward systematically classifying texts’ safety into categories, enabling more comprehensive safety evaluations. With the proliferation of new (AI) policies in numerous countries (EU (2023), UK (2023), or US (2023)), there is a pressing need for expansive and adaptable taxonomies across modalities. Recent works like LlamaGuard2 (Team, 2024b) and MLCommons (Vidgen et al., 2024) and AIR 2024 (Zeng et al., 2024) mark significant advancements in this direction.
@@
Concurrent Approaches for Moderating Images. Along these lines, various approaches have been investigated, leveraging advanced models. For instance, leveraging large multimodal models’ underlying capabilities and comprehensive understanding of the real world, can be employed for visual content moderation. While prominent tools such as GPT-4 (OpenAI, 2024a) or Gemini (Team, 2024a) often remain closed-source, several open-source alternatives including Llava (Liu et al., 2023a, b), InternVL (Chen et al., 2024) and QwenVL (Wang et al., 2024) are available. However, these models lack a safety-specific understanding. Therefore, recent studies have fine-tuned them for content moderation, including LlamaGuard-3-Vision (Chi et al., 2024), ImageGuard (Li et al., 2025), OpenAI’s Omni-Moderation (OpenAI, 2024b). LlamaGuard-3-Vision focuses on safeguarding Human-AI conversations, which is insufficient for moderating images, as we demonstrate. While OpenAI’s Omni Moderation and ImageGuard are developed for content moderation, they fail at performing well on the task in general and cannot handle policies flexibly. Furthermore, OpenAI’s Omni Moderation is closed-source.
In contrast, LlavaGuard is an open-framework that performs well in moderating visual content while offering the flexibility to adapt to different policies, making it the first robust tool available for this purpose.
@@ 3 LlavaGuard’s Safety Taxonomy for Vision
Creating automated safety checks for visual inputs requires classifiers to analyze and assess images in real time. A well-defined safety taxonomy is a foundational component for building such systems. To this end, we have developed a flexible taxonomy focused on safety categories and risk guidelines to identify and address unsafe image content. This taxonomy serves as a default framework that enables training on diverse policies and can, in turn, be easily adapted to various use cases by modifying, including extending or removing, the safety categories and the risk guidelines. A detailed overview of our safety taxonomy is available in App. 3.
@@
Safety Categories. A key element of our taxonomy is the set of safety categories outlined in Tab. 1. Distinct from existing safety taxonomies (Inan et al., 2023; Wang et al., 2023), our taxonomy is uniquely designed for the vision domain. It incorporates the latest AI regulations (EU, 2023; UK, 2023; US, 2023) and includes nine safety categories, along with an additional NA category for content that doesn’t pertain to any safety concerns and is therefore always considered safe (cf. Tab. 1). To this end, we have expanded upon existing text-based categories, introducing new distinctions and categories tailored to the visual domain. For example, Nudity and Sexual Content are more pertinent to visual content, while Disasters or Emergencies are newly included for assessing image safety. For future reference, we will use category shortcuts (e.g., O3 or NA).
Risk Guidelines. Each safety category is defined by a detailed description, i.e. risk guideline, to elicit an in-depth safety understanding. These guidelines specify what explicitly should not and what can be included. For example, without such a detailed guideline, the model might ban all forms of nudity, although it may remain important, e.g., for the educational and medical domains. Furthermore, this setup can flexibly adjust the safety policy to varying contexts and settings, e.g., by moving certain bullet points from Should not to Can and vice versa. Further, we may entirely disregard a certain category by using only one set of guidelines preceded by a statement like ‘Category O6 is declared as non-violating. Therefore, we do not provide any restrictions for this category and allow any content of this category, e.g. ...’. By providing explicit instructions outlining what is permitted and what is not, we achieve greater control over how the model adheres to a given safety policy in its evaluation.
@@ 4 Dataset Creation
Data Collection. We used the Socio-Moral Image Database (SMID) (Crone et al., 2018) as the foundation of our safety data collection. The SMID dataset is a human-created set of images annotated on various safety dimensions. While this dataset serves as a solid basis, it suffers from a large imbalance in the number of images per safety category. Specifically, most SMID images depict violence or hate while there are nearly none depicting sexual content and only a few self-harm or animal cruelty. To achieve a better balance among the categories, we extended the dataset with web-crawled images. To this end, we web-scraped images from Google and Bing Search, collecting enough images to ensure each category contains at least 100 images of different safety severity levels.
@@
Data Augmentation. A universal safeguard should be able to adapt its assessment to varying safety taxonomies. To promote this behavior, we implement two data augmentation techniques. First, we introduce additional samples with a modified policy prompt. Specifically, we pick samples initially rated as unsafe and declare the violated category as non-violating, thus flipping the respective safety rating from unsafe to safe. These modified samples are subsequently referred to as policy exceptions. Second, we add further samples where we declare up to 3 random safety categories as non-violating. These categories are selected so that the violated category remains untouched.
Guided Rationales. While ratings and categories are essential annotations for image safety, they offer only limited insight into the image content and the underlying rationale for safety assessments. To bridge this gap, we introduce rationales that clearly explain why an image is rated as safe or unsafe. Providing rationales comes with two decisive benefits: (1) They improve transparency by showing users the basis for each assessment, clarifying how specific image features relate to the assigned safety label. (2) They enable VLMs to learn the correct reasoning behind safety assessments and enhance the models’ interpretability. Accordingly, when rationales are used for training, their quality is of particular importance since high-quality rationales allow the VLMs to learn a more nuanced safety understanding. Unfortunately, collecting such detailed rationales is very difficult. Human annotation is time-consuming and prevents the dataset from being easily expanded. On the other side, naïve generation (based on policy and image) often yields incoherent rationales that fail to capture policy-specific nuances or even disregard the policy entirely (cf. App. Fig. 7(a)).
@@
Dataset Construction. Finally, we constructed a dataset comprising 5,466 unique (3,242 safe and 2,224 unsafe) samples, each annotated with a safety rating, category, and guided rationale. 3,242 samples are based on the default policy, while the remainder use augmented policies. The dataset is split into 4571 (train), 71 (eval), and 824 (test). The test set is balanced across safety categories and ratings. App. I provides additional insights into the composition of the dataset. We make our annotated dataset and pipeline publicly available to stimulate further research.

### v3

@@ Abstract
This paper introduces LlavaGuard, a suite of VLM-based vision safeguards that address the critical need for reliable guardrails in the era of large-scale data and models. To this end, we establish a novel open framework, describing a customizable safety taxonomy, data preprocessing, augmentation, and training setup. For teaching a VLM safeguard on safety, we further create a multimodal safety dataset with high-quality human expert annotations, where each image is labeled with a safety rating, category, and rationale. We also employ advanced augmentations to support context-specific assessments. The resulting LlavaGuard models, ranging from 0.5B to 7B, serve as a versatile tool for evaluating the safety compliance of visual content against flexible policies. In comprehensive experiments, LlavaGuard outperforms both state-of-the-art safeguards and VLMs in accuracy and in flexibly handling different policies. Additionally, we demonstrate LlavaGuard’s performance in two real-world applications: large-scale dataset annotation and moderation of text-to-image models. We make our entire framework, including the dataset, model weights, and training code, publicly available at https://ml-research.github.io/human-centered-genai/projects/llavaguard.
Content Moderation, AI Safety, Responsible AI, Safeguards
Warning: This paper contains explicit imagery and other content that readers may find disturbing.
@@ 1 Introduction
Furthermore, emerging legal frameworks for AI, such as those in the EU (EU, 2023), US (US, 2023), and UK (UK, 2023), are pushing for advanced generative models to comply with new regulations. This has led to the proposal of various safety approaches and taxonomies to systematically assess and mitigate the risks associated with large-scale data and models (Inan et al., 2023; Wang et al., 2023; Schramowski et al., 2023; Tedeschi et al., 2024). However, prior safety research focuses primarily on the text domain, leaving a distinct lack of frameworks for the visual modality. Moreover, there is a dearth of datasets and pipelines necessary for building advanced safeguards. Consequently, users largely have to rely on rigid NSFW classifications (Qu et al., 2024; Birhane & Prabhu, 2021; Birhane et al., 2023; Schramowski et al., 2022; NotAI-tech, 2019; Laborde, 2020), which lack the context-awareness and flexibility needed for more nuanced, fine-grained analysis.
We bridge this gap by introducing LlavaGuard (Fig. 1), a versatile framework for assessing potentially unsafe image content. LlavaGuard combines visual and textual inputs, allowing for the assessment of arbitrary safety policies to meet diverse requirements. To this end, we enhance the general capabilities of VLMs in two key ways. Firstly, building on their inherent common-sense understanding, LlavaGuard is trained with an in-depth and adaptive understanding of safety. This safety-specific training enables it to provide detailed responses that include an overall safety rating (safe/unsafe), a specific safety category (e.g. hate or sexual content), and a rationale that explains why the content is deemed unsafe according to the given policy. Secondly, with an advanced training setup, LlavaGuard is equipped with the ability to flexibly handle a broad spectrum of policies. Given the variability in regulations—such as cannabis being illegal in some countries but legal in others—LlavaGuard can be easily adjusted to both contexts.
In summary, our contributions are as follows: (1) We establish an open framework for vision safeguards, encompassing a safety taxonomy, data preprocessing, augmentation, and training setup. (2) We construct a multimodal safety dataset with human annotations, including images labeled with a safety rating, category, and rationale (Sec. 4). (3) Based on the previous, we launch LlavaGuard, a suite of vision safeguards based on VLMs, trained to assess visual content for safety (Sec. 5). (4) We conduct comprehensive experiments demonstrating that LlavaGuard outperforms state-of-the-art VLMs and state-of-the-art safeguards, excelling not only in accuracy but also in flexibly handling different policies (Sec. 6). (5) Finally, we validate LlavaGuard’s performance on two real-world applications: dataset annotation and moderation of generative models (Sec. 7).
@@ 2 Background
Gebru et al. (2021) initiated the effort of systematically reporting visual content by advocating for meticulous documentation of datasets to promote their ethical use. Initial approaches are centered around classification tools, where common ones are convolutional (NotAI-tech, 2019; Karkkainen & Joo, 2021) and CLIP-based (Schramowski et al., 2022; Nichol et al., 2022) classifiers or human annotations (Birhane et al., 2021). In particular, NudeNet (NotAI-tech, 2019), NSFW-Nets (Falconsai, 2024; Sanali209, 2024) and Birhane & Prabhu (2021) focus on NSFW, FairFace (Karkkainen & Joo, 2021) on fairness, Q16 (Schramowski et al., 2022) on in/appropriatness, and Nichol et al. (2022) on privacy and violence.
@@
Based on these tools and efforts, common large-scale datasets such as LAION (Schuhmann et al., 2022) or ImageNet (Deng et al., 2009) have undergone careful curation from different perspectives (Qu et al., 2024; Birhane & Prabhu, 2021; Birhane et al., 2023; Schramowski et al., 2022; Schuhmann et al., 2022). The resulting (unsafe) subsets serve a dual purpose. First, they are crucial in excluding content that could compromise safety during model training, ensuring a safer training environment. Second, these subsets provide valuable resources for conducting safety-oriented research. Furthermore, with the rise of models generating images, prompt testbeds such as I2P (Schramowski et al., 2023) or MAGBIG (Friedrich et al., 2024b) have been proposed for safety audits, moving beyond real images.
@@
However, the scope of these audits and auditing tools is limited by the capabilities of their underlying models, which lack the versatility and advanced common-sense understanding provided by large-scale, pre-trained VLMs. This capability is essential for effectively handling both real and synthetic images across a broad range of domains. In contrast, LlavaGuard is built on such VLMs, which are not restricted to a fixed set of safety dimensions and can be easily adapted to accommodate a variety of policies.
Generative AI Risk Assessment and Mitigation.
@@
In the context of generative models, most existing studies focus on the textual modality. Endeavors to systematically categorize safety risks have spurred the creation of safety taxonomies (Inan et al., 2023; Wang et al., 2023; Tedeschi et al., 2024), which provide a structured framework for assessing and mitigating risks. In particular, Inan et al. (2023) proposed a taxonomy enabling the LlamaGuard model to classify harmful prompts and responses into six categories. Similarly, Wang et al. (2023) proposed an 8-category taxonomy to evaluate LLMs based on different safety and trustworthiness perspectives, including robustness to adversarial attacks. These taxonomies constitute an initial stride toward systematically classifying texts’ safety into categories, enabling more comprehensive safety evaluations. With the proliferation of new (AI) policies in numerous countries (EU (2023), UK (2023), or US (2023)), there is a pressing need for expansive and adaptable taxonomies across modalities. Recent works like LlamaGuard2 (Team, 2024b), MLCommons (Vidgen et al., 2024), and AIR 2024 (Zeng et al., 2024) mark significant advancements in this direction.
@@
Concurrent Approaches for Moderating Images. Along these lines, various approaches have been investigated, leveraging advanced models. For instance, leveraging large multimodal models’ underlying capabilities and comprehensive understanding of the real world can be employed for visual content moderation. While prominent tools such as GPT-4 (OpenAI, 2024a) or Gemini (Team, 2024a) often remain closed-source, several open-source alternatives including Llava (Liu et al., 2023a, b), InternVL (Chen et al., 2024) and QwenVL (Wang et al., 2024) are available. However, these models lack a safety-specific understanding. Therefore, recent studies have fine-tuned them for content moderation, including LlamaGuard-3-Vision (Chi et al., 2024), ImageGuard (Li et al., 2025), and OpenAI’s Omni-Moderation (OpenAI, 2024b). LlamaGuard-3-Vision focuses on safeguarding human-AI conversations, which is insufficient for moderating images, as we demonstrate. While OpenAI’s Omni Moderation and ImageGuard are developed for content moderation, they fail at performing well on the task in general and cannot handle policies flexibly. Furthermore, OpenAI’s Omni Moderation is closed-source.
In contrast, LlavaGuard is an open framework that performs well in moderating visual content while offering the flexibility to adapt to different policies, making it the first robust tool available for this purpose.
@@ 3 LlavaGuard’s Safety Taxonomy for Vision
Creating automated safety checks for visual inputs requires classifiers to analyze and assess images in real-time. A well-defined safety taxonomy is a foundational component for building such systems. To this end, we have developed a flexible taxonomy focused on safety categories and risk guidelines to identify and address unsafe image content. This taxonomy serves as a default framework that enables training on diverse policies and can, in turn, be easily adapted to various use cases by modifying, including extending or removing, the safety categories and the risk guidelines. A detailed overview of our safety taxonomy is available in App. 3.
@@
3.1 Safety Categories
A key element of our taxonomy is the set of safety categories outlined in Tab. 1. Distinct from existing safety taxonomies (Inan et al., 2023; Wang et al., 2023), our taxonomy is uniquely designed for the vision domain. It incorporates the latest AI regulations (EU, 2023; UK, 2023; US, 2023) and includes nine safety categories, along with an additional NA category for content that doesn’t pertain to any safety concerns and is therefore always considered safe (cf. Tab. 1). To this end, we have expanded upon existing text-based categories, introducing new distinctions and categories tailored to the visual domain. For example, Nudity and Sexual Content are more pertinent to visual content, while Disasters or Emergencies are newly included for assessing image safety. For future reference, we will use category shortcuts (e.g., O3 or NA).
3.2 Risk Guidelines
Each safety category is defined by a detailed description, i.e. risk guideline, to elicit an in-depth safety understanding. These guidelines specify what explicitly should not and what can be included. For example, without such a detailed guideline, the model might ban all forms of nudity, although it may remain important, e.g., for the educational and medical domains. Furthermore, this setup can flexibly adjust the safety policy to varying contexts and settings, e.g., by moving certain bullet points from Should not to Can and vice versa. Further, we may entirely disregard a certain category by using only one set of guidelines preceded by a statement like ‘Category O6 is declared as non-violating. Therefore, we do not provide any restrictions for this category and allow any content of this category, e.g. ...’. By providing explicit instructions outlining what is permitted and what is not, we achieve greater control over how the model adheres to a given safety policy in its evaluation.
@@ 4 Dataset Creation
Table 2: Guided vs. non-guided rationales, judged by GPT-4o on a scale from 1 to 10. Top: Comparison of guided and base rationales for Llava-34B. Bottom: Guided rationale quality across model scales. Guided rationales are markedly superior, with Llava-34B performing best overall.
Model Type Mean Median Win (%)
Guided vs. Base
Llava-34B Base 3.8 3.0 0.1
Llava-34B Guided 9.1 9.0 99.9
Rationale Quality Across Model Sizes
Llava-7B Guided 7.0 6.8 6.6
Llava-13B Guided 7.0 6.7 9.6
Llava-34B Guided 9.0 8.4 83.9
4.1 Data Collection
We used the Socio-Moral Image Database (SMID) (Crone et al., 2018) as the foundation of our safety data collection. The SMID dataset is a human-created set of images annotated on various safety dimensions. While this dataset serves as a solid basis, it suffers from a large imbalance in the number of images per safety category. Specifically, most SMID images depict violence or hate while there are nearly none depicting sexual content and only a few self-harm or animal cruelty. To achieve a better balance among the categories, we extended the dataset with web-crawled images. To this end, we web-scraped images from Google and Bing Search, collecting enough images to ensure each category contains at least 100 images of different safety severity levels.
@@ 4.1 Data Collection
4.2 Data Augmentation
A universal safeguard should be able to adapt its assessment to varying safety taxonomies. To promote this behavior, we implement two data augmentation techniques. First, we introduce additional samples with a modified policy prompt. Specifically, we pick samples initially rated as unsafe and declare the violated category as non-violating, thus flipping the respective safety rating from unsafe to safe. These modified samples are subsequently referred to as policy exceptions. Second, we add further samples where we declare up to 3 random safety categories as non-violating. These categories are selected so that the violated category remains untouched.
4.3 Guided Rationales
While ratings and categories are essential annotations for image safety, they offer only limited insight into the image content and the underlying rationale for safety assessments. To bridge this gap, we introduce rationales that clearly explain why an image is rated as safe or unsafe. Providing rationales comes with two decisive benefits: (1) They improve transparency by showing users the basis for each assessment, clarifying how specific image features relate to the assigned safety label. (2) They enable VLMs to learn the correct reasoning behind safety assessments and enhance the models’ interpretability. Accordingly, when rationales are used for training, their quality is of particular importance since high-quality rationales allow the VLMs to learn a more nuanced safety understanding. Unfortunately, collecting such detailed rationales is very difficult. Human annotation is time-consuming and prevents the dataset from being easily expanded. On the other side, naïve generation (based on policy and image) often yields incoherent rationales that fail to capture policy-specific nuances or even disregard the policy entirely (cf. App. Fig. 7(a)).
@@ 4.3 Guided Rationales
To validate the quality of the guided rationales, we evaluate all dataset rationales using GPT-4o, which scored them for comprehensiveness, accuracy, and adherence to policy guidelines. Table 2 presents the comparative results comparing guided vs. non-guided rationales as well as rationales generated from different Llava model scales. Guided rationales generated with Llava-34B achieve a substantially higher mean quality score (9.1) and median score (9.0) than their non-guided counterparts (mean 3.8, median 3.0), with a win rate of 99.9%. Moreover, when benchmarking across model scales, Llava-34B demonstrates the highest overall performance (win rate 83.9%, mean 9.0, median 8.4), greatly surpassing Llava-13B and Llava-7B in both quality and consistency. Furthermore, App. Fig. 7 indicates that the improved quality of guided rationales during training effectively translates to high-quality rationales post-training.
4.4 Dataset Construction
Finally, we constructed a dataset comprising 5,466 unique (3,242 safe and 2,224 unsafe) samples, each annotated with a safety rating, category, and guided rationale. 3,242 samples are based on the default policy, while the remainder use augmented policies. The dataset is split into 4571 (train), 71 (eval), and 824 (test). The test set is balanced across safety categories and ratings. App. I provides additional insights into the composition of the dataset.